A 0.5kg egg is falling at a speed of 10 m/s before hitting the ground. What is the force of impact if the impact time is measured to be 3.0 milliseconds?

1
Expert's answer
2022-05-04T12:40:21-0400

The Newton's second law says

"F=ma=m\frac{\Delta v}{\Delta t}"

"F=\rm 0.5\: kg*\frac{10\: m/s}{3.0*10^{-3}\: s}=1.7*10^3\: N"